Two minor league baseball players got a total of 386 hits. Washington had 12 more hits than Sanchez. Find the number of hits for each player.
Answer by checkley77(12569)

W+S=386
W=S+12
(S+12)=S=386
2S=386-12
2S=374
S=374/2
S=187 HITS.
W=187+12=199 HITS.
PROOF:
187+199=386
386=386